Please help me solve this problem: Find two consecutive integers whose product is five less than the square of the smaller number. Thanks, Bill
Answer by checkley71(8403)

x(x+1)=x^-5
x^2+x=x^2-5
x=-5 answer.
proof
-5(-5+1)=-5^2-5
-5*-4=25-5
20=20